OpenStudy (anonymous):

find p(x>2) x = P(X) 1 =0.30 2 =0.40 3 =0.15 4=?

hartnn (hartnn):

so when x=4, u want P(X) =... ? right ? use the property that sum of P(1) through P(4) =1 then u get P(4) easily.

hartnn (hartnn):

then P(x>2) is just P(3)+P(4)

OpenStudy (anonymous):

can you just do it for me.. how do i find p(4)?

OpenStudy (anonymous):

can you just do it for me.. how do i find p(4)?

hartnn (hartnn):

sorry, i can't just do it for you. i can show you what to do P(1)+P(2)+P(3)+P(4) = 1 you can find P(4) from that.

OpenStudy (anonymous):

so i would do 0.30+0.40+0.15? =0.85..then what do i do

OpenStudy (anonymous):

i loks like it might be 0.20

OpenStudy (anonymous):

howd you get that though..

OpenStudy (phi):

I don't understand the question. We can find P(x>2) using the idea that P(all x) = 1 and P(x>2) = 1 - P(x ≤ 2) however, unless we know that x can only be 1,2,3,4 we cannot find P(x=4)

hartnn (hartnn):

P(4) = 1-0.85 = 0.15 P(X>2) = P(3)+ P(4) = 0.15+0.15 = 0.3